OpenStudy (anonymous):

QUal é o resultado de (-8)³

OpenStudy (anonymous):

\[-8\times-8\times-8=-512\]

OpenStudy (anonymous):

whenever you see an odd number of power, it is always negative

OpenStudy (anonymous):

u can also solve it\[(-8)^{3}=(-1)^{3} *(8)^{3}=(-1)*(512)=-512\] Try both one suggested by @ashleycheng123 and me. whichever u find easier carry on with it